Assume a major investment service has just given Oasis Electronics its highest investment​ rating, along with a strong buy recommendation. As a​ result, you decide to take a look for yourself and to place a value on the​ company's stock.​ Here's what you​ find: This​ year, Oasis paid its stockholders an annual dividend of $3.01 a​ share, but because of its high rate of growth in​ earnings, its dividends are expected to grow at the rate of 16​% a year for the next 4 years and then to level out at 13% a year. So​ far, you've learned that the stock has a beta of 1.95 the​ risk-free rate of return is 5​%,and the expected return on the market is 12​%.

Using the CAPM to find the required rate of​ return, put a value on this stock.
Answer #1

D0 = 3.01

D1 = 3.01( 1+ 0.16) = 3.4916

D2 = 3.4916( 1+ 0.16) = 4.050256

D3 = 4.050256( 1+ 0.16) = 4.69829696

D4 = 4.69829696( 1+ 0.16) = 5.4500244736

As per CAPM, Required Rate of Return

= 5% + 1.95( 12% - 5%)

= 18.65%

r is 18.65%

G is 13%

= $ 66.38